# Quantum Doubly Stochastic Transformers

### v1

@@ Abstract
At the core of the Transformer, the softmax normalizes the attention matrix to be right stochastic. Previous research has shown that this often de-stabilizes training and that enforcing the attention matrix to be doubly stochastic (through Sinkhorn’s algorithm) consistently improves performance across different tasks, domains and Transformer flavors. However, Sinkhorn’s algorithm is iterative, approximative, non-parametric and thus inflexible w.r.t. the obtained doubly stochastic matrix (DSM). Recently, it has been proven that DSMs can be obtained with a parametric quantum circuit, yielding a novel quantum inductive bias for DSMs with no known classical analogue. Motivated by this, we demonstrate the feasibility of a hybrid classical-quantum doubly stochastic Transformer (QDSFormer) that replaces the softmax in the self-attention layer with a variational quantum circuit. We study the expressive power of the circuit and find that it yields more diverse DSMs that better preserve information than classical operators. Across multiple small-scale object recognition tasks, we find that our QDSFormer consistently surpasses both a standard Vision Transformer and other doubly stochastic Transformers. Beyond the established Sinkformer, this comparison includes a novel quantum-inspired doubly stochastic Transformer (based on QR decomposition) that can be of independent interest. The QDSFormer also shows improved training stability and lower performance variation suggesting that it may mitigate the notoriously unstable training of ViTs on small-scale data.
@@ 1 Introduction
The Transformer [52] continues to be a dominant building block powering models in natural language processing [17], computer vision [16, 32], chemistry [8] and biology [3]. At the same time, quantum computing (QC) is a novel paradigm with the potential to become practically useful in ML [43, 23, 34, 25, 1] and fuel applications across disciplines [5, 14]. Many attempts have been made to build Transformers with quantum gates, either entirely [29, 39] or only the attention blocks [28, 19, 51]. However, rather than merely migrating, recent work in quantum ML identified constraints in specific flavors of neural networks (NN) and successfully mitigated those through quantum – e.g., fourier NNs [63], graph NNs [50] or input-convex NNs [37]. Some known limitations of Transformers are due to the softmax in the attention block, e.g., entropy collapse [62], rank collapse [40], token uniformity [15], eureka moments [24] and more [59, 11, 45, 12, 61]. Applying softmax enforces the attention matrices to be right-stochastic (i.e., rows sum to $1$). Its temperature controls the entropy of this distribution, which is often adjusted to stabilize training [24, 40].
@@
Concurrently, Sander et al. [42] discovered that Transformer attention naturally converge to doubly stochastic matrices (DSMs) over training, i.e. their rows and columns sum to 1. Motivated by this, they discovered that enforcing bistochasticity boosts Transformer performance across different modalities (text, images, point clouds). Intuitively, doubly stochastic attention has a similar effect to increasing temperature (entropy) – attention becomes more ”democratic”, less interactions are missed and all tokens are being attended more equally. Their Sinkformer [42] is a generalization of Transformers that leverages Sinkhorn’s algorithm (SA) and has been widely adopted and extended [21, 60, 44]. Among various techniques to obtain DSMs [46, 41, 13, 54], SA is the most obvious choice, however it has some disadvantages. (1) It is an iterative approximation procedure which reaches a DSM only in the limit. It is thus empirical how many iterations a Sinkformer needs. (2) It can guarantee to find a DSMs only if the input matrix is non-negative, which is generally not the case within a Transformer. (3) It is non-parametric. Thus, in contrast to e.g., a NN layer, it cannot be optimized regarding which DSM should be returned. Given the empirical superiority of the Sinkformer to vanilla Transformers, it is natural to study different techniques to make attention doubly stochastic.
@@
Strikingly, the recent work by Mariella et al. [37] proved (in a different context) that DSMs can be obtained naturally with a variational (i.e., parametric) quantum circuit, dubbed QontOT. They emphasize that there exists no classical learning (i.e., parametric) method that can produce a DSM, akin to QontOT. Here, we demonstrate that this opens the door for a hybrid quantum-classical doubly-stochastic Transformer (QDSFormer) which offers more flexibility than the Sinkformer. To that end, we extend QontOT to emit DSMs for an equally-sized matrix. The resulting quantum layer may replace the softmax within any standard (i.e., non-local, non-sparse) self-attention block. We focus on replacing the softmax inside the scaled dot-product self-attention of a Vision Transformer (ViT) for three reasons:
@@
    ViTs [16] suffer from unstable training [62, 24]
@@
    Unlike in NLP, the attention matrix has constant size, which eases the application of the quantum circuit
@@
    The attention matrix in a Transformer encoder is unconstrained (decoder attention matrices are masked).
We empirically analyse expressivity of the quantum circuit, finding that it yields more diverse DSMs than Sinkhorn’s algorithm both on synthetic and real data. It also preserves information better and induces higher entropy. We then train various flavors of doubly stochastic Transformers (see Figure 1) on more than ten object recognition datasets. In comparison to the ViT [16] and Sinkformer [42], the QDSFormer shows competitive performance, consistently surpassing both while stabilizing training.
@@ 2.1 Doubly Stochastic Matrices (DSMs)
We denote the $n$-dimensional vector of ones by $\bm{\mathrm{1}}_{n}$ and the $n\times n$ identity matrix as $\bm{\mathrm{I}}_{n}$. The Birkhoff polytope $\Omega_{n}\coloneqq\mathcal{N}\left(\bm{\mathrm{1}}_{n},\bm{\mathrm{1}}_{n}\right)$ [9] defines the convex set of $n\times n$ doubly stochastic matrices (DSMs). A DSM $\bm{\mathrm{P}}\in\Omega_{n}$ is a non-negative matrix with row/column sum of 1
@@
A right stochastic matrix $\bm{\mathrm{R}}$ has row sums of 1
$\bm{\mathrm{R}}\bm{\mathrm{1}}_{n}=\bm{\mathrm{1}}_{n},\quad\bm{\mathrm{R}}_{i,j}\geq 0,$ (2)
and a left stochastic matrix $\bm{\mathrm{L}}$ has column sums of 1
$\bm{\mathrm{L}}^{\top}\bm{\mathrm{1}}_{n}=\bm{\mathrm{1}}_{n},\quad\bm{\mathrm{L}}_{i,j}\geq 0.$ (3)
Hence, a DSM is left and right stochastic. Moreover, the Birkhoff-von Neumann theorem states that the $n!$ vertices (i.e., extreme points) of the Birkhoff polytope $\Omega_{n}$ are permutation matrices, so their entries belong to $\{0,1\}$. Notably, every DSM $\bm{\mathrm{P}}\in\Omega_{n}$ can be decomposed as a convex combination of permutation matrices, that is $\bm{\mathrm{P}}=\sum_{i=1}^{N}\lambda_{i}\bm{\mathrm{\Pi}}_{i}$ for some probability vector $\bm{\lambda}\in\Delta_{N}$, $n\times n$ permutation matrices $\{\bm{\mathrm{\Pi}}_{i}\}$, and the number of extreme points $N\leq n^{2}$. While the decomposition is not unique, each DSM can be represented by at most $n^{2}$ permutation matrices [7]. Due to the linear equality constraints, the Birkhoff polytope $\Omega_{n}$ lies within a $(n-1)^{2}$-dimensional affine subspace of the $n^{2}$ dimensional space of $\mathbb{R}^{n\times n}$ matrices.
@@ 2.2 Attention
We study extensions of dot product attention [52]
@@
where $\bm{\mathrm{Q}}\coloneqq\bm{\mathrm{X}}\bm{\mathrm{W}}_{Q}$, $\bm{\mathrm{K}}\coloneqq\bm{\mathrm{X}}\bm{\mathrm{W}}_{K}$ and $\bm{\mathrm{V}}\coloneqq\bm{\mathrm{X}}\bm{\mathrm{W}}_{V}$ map the input $\bm{\mathrm{X}}$ to query $\bm{\mathrm{Q}}$, key $\bm{\mathrm{K}}$ and value $\bm{\mathrm{V}}$ through their respective weight matrix $\bm{\mathrm{W}}_{i}$ s.t. $\bm{\mathrm{Q}},\bm{\mathrm{K}},\bm{\mathrm{V}}\in\mathbb{R}^{T\times d}$. Note that $\tau^{-1}\bm{\mathrm{Q}}\bm{\mathrm{K}}^{\top}\in\mathbb{R}^{T\times T}$, so the unnormalized attention matrix is quadratic. Applying the softmax operator, denoted $S(\bm{\mathrm{z}})_{i}=\frac{\exp(\bm{\mathrm{z}}_{i})}{\sum_{j=1}^{n}\exp(\bm{\mathrm{z}}_{j})}$, over the rows makes $\bm{\mathrm{A}}$ right-stochastic, i.e., each row $i$ contains a probability distribution denoting the amount of “attention” token $i$ pays to the other tokens. The temperature $\tau$, canonically set to $\sqrt{d_{k}}$ [52], controls the entropy of the output: low temperature yields a peaky distribution emphasizing differences whereas high temperature attenuates differences thus increasing entropy.
@@ 2.3 Doubly-Stochastic Operators
Below we define a non-exhaustive set of operators that can transform $\bm{\mathrm{M}}\in\mathbb{R_{+}}^{T\times T}$ to a DSM $\bm{\mathrm{P}}\in\Omega_{T}$. The operators can be integrated into a Transformer by $\bm{\mathrm{M}}\coloneqq\bm{\mathrm{Q}}\bm{\mathrm{K}}^{\top}$ thus yielding a Doubly Stochastic Transformer (”DSFormer”).
@@ 2.3.1 Sinkhorn’s algorithm
The most natural approach to obtain a doubly stochastic Transformer was pursued in the Sinkformer [42] and leverages Sinkhorn’s algorithm [46]. Sinkhorn’s algorithm is based on Sinkhorn’s theorem, stating that for any square strictly positive matrix $\bm{\mathrm{M}}\in\mathbb{R_{+}}^{T\times T}$, there exist (strictly) positive diagonal matrices $\bm{\mathrm{P}}=\bm{\mathrm{D}}_{1}$, $\bm{\mathrm{D}}_{2}$ s.t., $\bm{\mathrm{D}}_{1}\bm{\mathrm{M}}\bm{\mathrm{D}}_{2}\in\Omega_{T}$. Sinkhorn’s algorithmis an approximation procedure that iteratively normalizes the mass of the rows and the columns of $\bm{\mathrm{M}}$ which has been proven to converge to a DSM by minimizing KL divergence [48]. The sole hyperparameter of this procedure is $K$, the number of iterations, which we enforce to be odd, following [42], to ensure the resulting matrix is at least numerically row-stochastic, like for the canonical Softmax operator. Moreover, we study and compare two implementations of Sinkhorn’s algorithm (SA), Naive and OT. Naive alternates between column- and row-normalization: at even iterations ($t$), each column is normalized as $\bm{\mathrm{P}}_{ij}^{(t+1)}=\bm{\mathrm{P}}_{ij}^{(t)}/{\sum_{i}\bm{\mathrm{P}}_{ij}^{(t)}}$, and at odd iterations, each row is normalized as $\bm{\mathrm{P}}_{ij}^{(t+1)}=\bm{\mathrm{P}}_{ij}^{(t)}/{\sum_{j}\bm{\mathrm{P}}_{ij}^{(t)}}$. Instead, the OT flavor is the operator used in the Sinkformer [42] which relies on the more robust and generalized version proposed by Cuturi [13] to compute optimal transport distances. Note that, both flavors may not converge with few iterations, especially if $\bm{\mathrm{Q}}\bm{\mathrm{K}}^{\top}$ contains large numeric values. Therefore, the Sinkformer is only an approximately doubly stochastic Transformer.

### v2

@@ Abstract
At the core of the Transformer, the softmax normalizes the attention matrix to be right stochastic. Previous research has shown that this often de-stabilizes training and that enforcing the attention matrix to be doubly stochastic (through Sinkhorn’s algorithm) consistently improves performance across different tasks, domains and Transformer flavors. However, Sinkhorn’s algorithm is iterative, approximative, non-parametric and thus inflexible w.r.t. the obtained doubly stochastic matrix (DSM). Recently, it has been proven that DSMs can be obtained with a parametric quantum circuit, yielding a novel quantum inductive bias for DSMs with no known classical analogue. Motivated by this, we demonstrate the feasibility of a hybrid classical-quantum doubly stochastic Transformer (QDSFormer) that replaces the softmax in the self-attention layer with a variational quantum circuit. We study the expressive power of the circuit and find that it yields more diverse DSMs that better preserve information than classical operators. Across multiple small-scale object recognition tasks, we find that our QDSFormer consistently surpasses both a standard ViT and other doubly stochastic Transformers. Beyond the Sinkformer, this comparison includes a novel quantum-inspired doubly stochastic Transformer (based on QR decomposition) that can be of independent interest. Our QDSFormer also shows improved training stability and lower performance variation suggesting that it may mitigate the notoriously unstable training of ViTs on small-scale data.
@@ 1 Introduction
The Transformer [vaswani2017attention] continues to be a dominant building block in natural language processing [dubey2024llama], computer vision [dosovitskiy2021an, kirillov2023segment] and biology [abramson2024accurate]. Quantum computing (QC), instead, is a novel paradigm with the potential to become practically useful in ML [schuld2022quantum, havlivcek2019supervised, Liu_2021QuantumSpeedUpML, Huang21_powerofdatainQML, abbas2021power] and fuel applications across disciplines [basu2023towards, di2024quantum]. Many attempts have been made to build Transformers with quantum gates, either entirely [khatri2024quixer, nguyen2024qclusformer, guo2024quantum] or only the attention blocks [kerenidis2024quantum, evans2024learning, unlu2024hybrid]. However, rather than merely migrating, recent work in quantum ML identified constraints in specific flavors of neural networks (NN) and successfully mitigated those through quantum – e.g., fourier NNs [zhao2024quantum], graph NNs [thabet2024quantum] or input-convex NNs [mariella2024quantum]. Some known limitations of Transformers are due to the softmax in the attention block, e.g., entropy collapse [zhai2023stabilizing], rank collapse [noci2022signal], token uniformity [dong2021attention], eureka moments [hoffmann2024eureka] and more [yang2019mixtape, chang2022softmax, shen2023study, chen2023accumulated, yuan2024towards]. Applying softmax enforces the attention matrices to be right-stochastic (i.e., rows sum to $1$) while its temperature controls the distribution entropy and is often adjusted to stabilize training [hoffmann2024eureka, noci2022signal].
@@
Concurrently, it was discovered that Transformer attention naturally converge to doubly stochastic matrices (DSMs) over training, i.e. their rows and columns sum to 1 [sander2022sinkformers]. Motivated by this, the Sinkformer [sander2022sinkformers] enforces bistochasticity which boosts Transformer performance across different modalities (text, images, point clouds). Intuitively, doubly stochastic attention has a similar effect to increasing temperature (entropy) – attention becomes more "democratic", less interactions are missed and all tokens are being attended more equally. The Sinkformer [sander2022sinkformers] is a generalization of Transformers that leverages Sinkhorn’s algorithm (SA) and has been widely adopted and extended [geshkovski2024emergence, ye2024otseg, shahbazi2025espformer]. Among various techniques to obtain DSMs [sinkhorn1964relationship, rontsis2020optimal, cuturi2013sinkhorn, wang2010learning], SA is the most obvious choice, however it has some disadvantages:
  1. 1.
    It is an iterative approximation procedure which reaches a DSM only in the limit. It is thus empirical how many iterations a Sinkformer needs and poor initialization can drastically deteriorate performance [thornton2023rethinking].
  2. 2.
    It can guarantee to find a DSMs only if the input matrix is non-negative, which is generally not the case within a Transformer (in practice non-negativity is enforced via exponentiation but we show that this hampers expressivity).
  3. 3.
    Backpropagating through SA often yields ill-conditioned and exploding/vanishing gradients when $\varepsilon$ is small. In practice, under early stopping, SA is a sublinearly convergent mirror-descent fixed-point solver rather than a simple, well-conditioned layer [leger2021gradient].
  4. 4.
    It is non-parametric. Thus, in contrast to e.g., a NN layer, it cannot be optimized regarding which DSM should be returned.
Given the empirical superiority of the Sinkformer to vanilla Transformers, it is natural to study different techniques to make attention doubly stochastic.
@@
Strikingly, it was recently proven (in a different context) that DSMs can be obtained naturally with a variational (i.e., parametric) quantum circuit, dubbed QontOT [mariella2024quantum]. They emphasize that there exists no classical learning (i.e., parametric) method that can produce a DSM, akin to QontOT. Here, we demonstrate that this opens the door for a hybrid quantum-classical doubly-stochastic Transformer (QDSFormer) which offers more flexibility than the Sinkformer. To that end, we extend QontOT to emit DSMs for an equally-sized matrix. The resulting quantum layer may replace the softmax within any standard (i.e., non-local, non-sparse) self-attention block. We focus on replacing the softmax inside the scaled dot-product self-attention of a Vision Transformer (ViT) for three reasons:
@@
    ViTs [dosovitskiy2021an] suffer from unstable training [zhai2023stabilizing, hoffmann2024eureka]
@@
    Unlike in NLP, the attention matrix size is constant, which eases quantum circuit application
@@
    The attention matrix in a Transformer encoder is unconstrained (unlike in decoders)
We empirically analyse expressivity of the quantum circuit, finding that it yields more diverse DSMs than Sinkhorn’s algorithm both on synthetic and real data. It also preserves information better and induces higher entropy. We then train various flavors of doubly stochastic Transformers (see Figure 1) on more than ten object recognition datasets. In comparison to the ViT [dosovitskiy2021an] and Sinkformer [sander2022sinkformers], the QDSFormer shows competitive performance, consistently surpassing both. In a compositional image recognition task [hoffmann2024eureka], we find that they stabilize Transformer training and accelerate learning as they antedate the Eureka moment in compositional problem solving.
In concurrent work, Shahbazi et al. have proposed the EPSFormer [shahbazi2025espformer] and the LOTFormer [shahbazi2025lotformer], two doubly stochastic Transformers that, just like our QDSFormer, overcomes the dependence on Sinkhorn’s algorithm to reach doubly stochastic attention. The ESPFormer [shahbazi2025espformer] achieves this with sliced OT which is faster than SA but still slower than standard attention. Their improvement, the LOTFormer [shahbazi2025lotformer] marries doubly stochastic and linear attention via conditional OT, yielding better performance and scaling than softmax attention. Due to their concurrent nature, a performance comparison to ESPFormer and LOTFormer is not included in this work.
@@ 2.1 Doubly Stochastic Matrices (DSMs)
We denote the $n$-dimensional vector of ones by $\bm{\mathrm{1}}_{n}$ and the $n\times n$ identity matrix as $\bm{\mathrm{I}}_{n}$. The Birkhoff polytope $\Omega_{n}\coloneqq\mathcal{N}\left(\bm{\mathrm{1}}_{n},\bm{\mathrm{1}}_{n}\right)$ [brualdi_2006] defines the convex set of $n\times n$ doubly stochastic matrices (DSMs). A DSM $\bm{\mathrm{P}}\in\Omega_{n}$ is a non-negative matrix with row/column sum of 1, i.e.,
@@
A right stochastic matrix $\bm{\mathrm{R}}$ has row sums of 1, i.e., $\bm{\mathrm{R}}\bm{\mathrm{1}}_{n}=\bm{\mathrm{1}}_{n},\quad\bm{\mathrm{R}}_{i,j}\geq 0,$ and a left stochastic matrix $\bm{\mathrm{L}}$ has column sums of 1, i.e., $\bm{\mathrm{L}}^{\top}\bm{\mathrm{1}}_{n}=\bm{\mathrm{1}}_{n},\quad\bm{\mathrm{L}}_{i,j}\geq 0.$ Hence, a DSM is left and right stochastic. Moreover, the Birkhoff-von Neumann theorem states that the $n!$ vertices (i.e., extreme points) of the Birkhoff polytope $\Omega_{n}$ are permutation matrices, so their entries belong to $\{0,1\}$. Notably, every DSM $\bm{\mathrm{P}}\in\Omega_{n}$ can be decomposed as a convex combination of permutation matrices: $\bm{\mathrm{P}}=\sum_{i=1}^{N}\lambda_{i}\bm{\mathrm{\Pi}}_{i}$. Here $\bm{\lambda}\in\Delta_{N}$ is some probability vector in the probability simplex (denoted as $\Delta_{N}$), $\{\bm{\mathrm{\Pi}}_{i}\}$ are the $n\times n$ permutation matrices and $N\leq n^{2}$ denotes the extreme points. While the decomposition is not unique, each DSM can be represented by at most $n^{2}$ permutation matrices [birkhoff1946tres]. Due to the linear equality constraints, the Birkhoff polytope $\Omega_{n}$ lies within a $(n-1)^{2}$-dimensional affine subspace of the space of $\mathbb{R}^{n\times n}$ matrices.
@@ 2.2 Attention
We study extensions of dot product attention [vaswani2017attention]
@@
where $\bm{\mathrm{Q}}\coloneqq\bm{\mathrm{X}}\bm{\mathrm{W}}_{Q}$, $\bm{\mathrm{K}}\coloneqq\bm{\mathrm{X}}\bm{\mathrm{W}}_{K}$ and $\bm{\mathrm{V}}\coloneqq\bm{\mathrm{X}}\bm{\mathrm{W}}_{V}$ map the input $\bm{\mathrm{X}}$ to query $\bm{\mathrm{Q}}$, key $\bm{\mathrm{K}}$ and value $\bm{\mathrm{V}}$ through their respective weight matrix $\bm{\mathrm{W}}_{i}$ s.t. $\bm{\mathrm{Q}},\bm{\mathrm{K}},\bm{\mathrm{V}}\in\mathbb{R}^{T\times d}$. Moreover, $\tau$ is called the "temperature" and canonically set to $\sqrt{d_{k}}$ [vaswani2017attention]. It controls the entropy of the output: low temperature yields a peaky distribution emphasizing differences. High temperature attenuates differences thus increasing entropy. Note that $\tau^{-1}\bm{\mathrm{Q}}\bm{\mathrm{K}}^{\top}\in\mathbb{R}^{T\times T}$, so the unnormalized attention matrix is quadratic. Applying the softmax operator, denoted $S(\bm{\mathrm{z}})_{i}=\frac{\exp(\bm{\mathrm{z}}_{i})}{\sum_{j=1}^{n}\exp(\bm{\mathrm{z}}_{j})}$, over the rows makes $\bm{\mathrm{A}}$ right-stochastic, i.e., each row $i$ contains a probability distribution denoting the amount of “attention” token $i$ pays to the other tokens. The temperature $\tau$,
@@ 2.3 Doubly-Stochastic Operators
Below we define a non-exhaustive set of operators that can transform $\bm{\mathrm{M}}\in\mathbb{R_{+}}^{T\times T}$ to a DSM $\bm{\mathrm{P}}\in\Omega_{T}$. The operators can be integrated into a Transformer by $\bm{\mathrm{M}}\coloneqq\bm{\mathrm{Q}}\bm{\mathrm{K}}^{\top}$ thus yielding a Doubly Stochastic Transformer ("DSFormer").
@@ 2.3.1 Sinkhorn’s algorithm
The most natural approach to obtain a doubly stochastic Transformer was pursued in the Sinkformer [sander2022sinkformers] and leverages Sinkhorn’s algorithm [sinkhorn1964relationship]. Sinkhorn’s algorithm is based on Sinkhorn’s theorem, stating that for any square strictly positive matrix $\bm{\mathrm{M}}\in\mathbb{R_{+}}^{T\times T}$, there exist (strictly) positive diagonal matrices $\bm{\mathrm{P}}=\bm{\mathrm{D}}_{1}$, $\bm{\mathrm{D}}_{2}$ s.t., $\bm{\mathrm{D}}_{1}\bm{\mathrm{M}}\bm{\mathrm{D}}_{2}\in\Omega_{T}$. Sinkhorn’s algorithm, also known as iterative proportional fitting [bacharach1965estimating], is an approximation procedure that iteratively normalizes the mass of the rows and the columns of $\bm{\mathrm{M}}$ which has been proven to converge to a DSM by minimizing Kullback-Leibler (KL) divergence [soules1991rate]. The sole hyperparameter of this procedure is $K$, the number of iterations, which we enforce to be odd, following [sander2022sinkformers], to ensure the resulting matrix is at least numerically row-stochastic, like for the canonical Softmax operator. Moreover, we study and compare two implementations of Sinkhorn’s algorithm (SA), Naive and OT. Naive alternates between column- and row-normalization: at even iterations ($t$), each column is normalized as $\bm{\mathrm{P}}_{ij}^{(t+1)}=\bm{\mathrm{P}}_{ij}^{(t)}/{\sum_{i}\bm{\mathrm{P}}_{ij}^{(t)}}$, and at odd iterations, each row is normalized as $\bm{\mathrm{P}}_{ij}^{(t+1)}=\bm{\mathrm{P}}_{ij}^{(t)}/{\sum_{j}\bm{\mathrm{P}}_{ij}^{(t)}}$. Instead, the OT flavor is the operator used in the Sinkformer [sander2022sinkformers] which relies on the more robust and generalized version to compute optimal transport distances [cuturi2013sinkhorn]. Note that, both flavors may not converge with few iterations, especially if $\bm{\mathrm{Q}}\bm{\mathrm{K}}^{\top}$ contains large numeric values. Therefore, the Sinkformer is only an approximately doubly stochastic Transformer.